What is the pH of a 0.179 M monoprotic acid whose Ka is 5.887 × 10−3?

pH =
Answer #1

Dissociation equilibrium of monoprotic acid(HA) is

HA(aq) + H2O(l) -------> A-(aq) + H3O+(aq)

Ka = [A-][H3O+] / [HA] = 5.887 ×10-3

Initial concentration

[HA] = 0.179

[A-] = 0

[H3O+] = 0

change in concentration

[HA] = - x

[A-] = + x

[HA] = + x

equilibrium concentration

[HA] = 0.179 - x

[A-] = x

[HA] = x

so,

x2/ ( 0.179 - x) = 5.887 ×10-3

solving for x

x = 0.02965

[H3O+] = 0.02965M

pH = -log[H3O+]

pH = - log( 0.02965M)

pH = 1.53